# Denoising and Reconstruction of Nonlinear Dynamics using Truncated Reservoir Computing

### v1

@@ Abstract
Measurements acquired from distributed physical systems are often sparse and noisy. Therefore, signal processing and system identification tools are required to mitigate noise effects and reconstruct unobserved dynamics from limited sensor data. However, this process is particularly challenging because the fundamental equations governing the dynamics are largely unavailable in practice. Reservoir Computing (RC) techniques have shown promise in efficiently simulating dynamical systems through an unstructured and efficient computation graph comprising a set of neurons with random connectivity. However, the potential of RC to operate in noisy regimes and distinguish noise from the primary dynamics of the system has not been fully explored. This paper presents a novel RC method for noise filtering and reconstructing nonlinear dynamics, offering a novel learning protocol associated with hyperparameter optimization. The performance of the RC in terms of noise intensity, noise frequency content, and drastic shifts in dynamical parameters are studied in two illustrative examples involving the nonlinear dynamics of the Lorenz attractor and adaptive exponential integrate-and-fire system (AdEx). It is shown that the denoising performance improves via truncating redundant nodes and edges of the computing reservoir, as well as properly optimizing the hyperparameters, e.g., the leakage rate, the spectral radius, the input connectivity, and the ridge regression parameter. Furthermore, the presented framework shows good generalization behavior when tested for reconstructing unseen attractors from the bifurcation diagram. Compared to the Extended Kalman Filter (EKF), the presented RC framework yields competitive accuracy at low signal-to-noise ratios (SNRs) and high-frequency ranges.
@@
Reconstruction of unobserved dynamical responses from sparse and noisy observations is a crucial and challenging task when dealing with unknown nonlinear systems. Filtering techniques have commonly been employed to handle such problems, but they often demand a thorough understanding of the underlying physics, such as the governing equations in terms of a physics-based state-space model. This paper presents a novel truncated computing reservoir to simulate nonlinear dynamics and implicitly recover the underlying physics from noisy measurements. It is hypothesized that reservoir computers can distinguish noise from dynamics up to a large extent and reconstruct unmeasured dynamical responses from noisy ones without having to involve any supplementary noise reduction techniques. Although the RC acts as a model-free mathematical modeling technique, which does not employ any state-space modeling technique, it offers competitive accuracy compared to the conventional filtering techniques at certain noise levels.
@@ I Introduction
Measurements of physical quantities from distributed systems often appear to be noisy and incomplete in covering all response quantities. This issue has inspired the development of a wide range of methods, which fall under two major classes, namely model-free and model-based methods, depending on whether they demand the knowledge of underlying physics. A good example of the former is wavelet-based denoising Donoho (1995), which does not require incorporating physical knowledge explicitly, whereas stochastic filtering techniques, including Kalman-type filters Anderson and Moore (1979), often demand a general understanding of the physics, characterized via state-space models. However, the foundation of both methods primarily rests upon a key hypothesis, implying that low and high-dimensional subsystems have distinguishable features in a specific latent space.
@@
Reservoir computing is a computational framework rooted in the broader domain of recurrent neural networks (RNNs) and dynamical systems theory. It leverages the dynamics of a fixed, randomly initialized high-dimensional system, known as the "reservoir," to transform input data into a hidden space representation. In this paradigm, only the output layer is trained through ridge regression Ghani et al. (2010), linearly combining the reservoir states to minimize the prediction errors. This significantly reduces the computational complexity associated with the training, mainly because only a system of linear equations needs to be solved coupled with a proper regularization. RC has been particularly effective in time-series processing Thorne et al. (2022); Shahi, Fenton, and Cherry (2022), speech recognition Ghani et al. (2010); Araujo et al. (2020), and chaotic system modeling Pathak et al. (2017); Krishnagopal et al. (2020); Carroll (2022) due to its ability to exploit the reservoir’s rich temporal dynamics and memory capacity. Heuristics of building the reservoir formulation, such as including an output feedback or a delayed input, have been shown to improve the results Jaurigue et al. (2021). The generalization behavior of RC has recently been studied, where a trained reservoir is examined for making predictions on unexplored attractors Flynn, Tsachouridis, and Amann (2021); Parlitz (2024). However, the optimal size and structure of the reservoir remain an open question since the reservoir networks are mostly initialized considering random connectivity. Only lately researchers have started working on structure-function relationships to discover network properties that are beneficial for a given learning task in the RC framework Yadav, Sinha, and Stender (2025); Yadav et al. (2025).
@@
RC has recently also been used as a model-free mathematical tool for signal-denoising and filtering tasks. Semenova et al. have studied the propagation of noise from the input nodes toward the reservoir outputs in recurrent and multilayer networks Semenova et al. (2019). Estébanez et al. have shown that, while noise can degrade the performance of RC-based analog computations, its introduction in the training phase can lead to improved robustness and accuracy Estébanez, Fischer, and Soriano (2019). Röhm et al. have shown that an RC trained based on a single noisy trajectory may still be able to reproduce unexplored attractors Röhm, Gauthier, and Fischer (2021). Nathe et al. have investigated the impact of measurement noise on the performance of RCs when used for reconstructing chaotic systems Nathe et al. (2023). Lee et al. have employed RC along with a nonlinear forward operator for efficient image denoising Lee et al. (2025). Choi and Kim have separated the signal from noise as the predictable part modeled by RC and acquired information about the additive or multiplicative nature of the noise, as well as noise statistical properties Choi and Kim (2024). Despite these advances in RC-based denoising approaches, a thorough investigation is still required on the potential of RC for discarding different noise processes and factors affecting denoising performance.
@@
This work proposes a novel learning protocol for denoising and reconstruction of nonlinear dynamics based on a truncated RC framework. For this purpose, we first train the reservoir on a specific data set using the conventional ridge regression approach and then apply a novel optimization approach to fine-tune the reservoir hyperparameters, including the ridge regression coefficient, the leakage rate, and the spectral density. Once the optimized reservoir is obtained, we truncate redundant nodes and edges according to the reconstruction accuracy of the validation data set using a novel pruning strategy. Two numerical examples are provided to demonstrate the proposed protocol and study the impact of noise intensity and frequency content on the performance of the RC. The generalization behavior of the proposed networks is demonstrated against drastic changes in the dynamical parameters and SNR. Comparison with the EKF is provided for the Lorenz attractor example, considering different noise levels. The method is further demonstrated using the non-smooth dynamics of the AdEx model.
@@
The remainder of this paper is organized as follows: first, the mathematical foundation of the RC framework will be reviewed, presenting the conventional ridge regression method. Then, the reservoir’s hyperparameters will be identified, for which an optimization approach will be presented. Truncation of the computing reservoir through a pruning approach will be discussed, and further comparisons will be made between the proposed RC framework and conventional filtering techniques from a theoretical standpoint. The presented approaches will next be demonstrated in two nonlinear dynamics problems, and finally, further discussions and conclusions will be drawn about the performance of the proposed truncated RC framework.
@@ II.1 Echo State Network
Several neural network architectures are available for constructing reservoir computers, among which Echo State Network (ESN) Jaeger (2007) and Liquid State Machines (LSM) Maass, Natschlaeger, and Markram (2002); Hazan and Manevitz (2012) are perhaps the most widely used ones. In this paper, ESN is employed for creating a reservoir whose formulation is described in terms of the latent dynamical states:
$\mathbf{r}(t_{i+1})=(1-{\it{\alpha}})\mathbf{r}(t_{i})+{\it{\alpha}}\ f(\mathbf{W}_{res}\mathbf{r}(t_{i})+\mathbf{W}_{in}\mathbf{u}(t_{i})+\mathbf{b})$ (1)
where $f(.)$ is the activation function, considered $\tanh(.)$, $\mathbf{r}(t_{i})$ is the reservoir state at time $t_{i}$, $\alpha$ is the leakage rate representing how much of the preceding state contributes to the present state, $\mathbf{W}_{res}$ is the reservoir weight matrix, $\mathbf{W}_{in}$ is the input weight matrix, $\mathbf{u}(t_{i})$ is the input vector at time $t_{i}$, and $\mathbf{b}$ is the bias term. Notably, an important parameter of the reservoir matrix ($\mathbf{W}_{res}$) is its spectral radius, denoted by $\gamma$ and calculated as the greatest eigenvalue of the reservoir matrix Ren and Ma (2022).
Given the above equation, a recursive formulation is available for calculating the reservoir states. By collecting all states into a new vector like $\mathbf{R}_{n}=[\mathbf{r}^{T}(t_{1}),...,\mathbf{r}^{T}(t_{n})]^{T}$, measurements of the physical system at different time steps comprised in $\mathbb{Y}_{n}=[\mathbf{y}^{T}(t_{1}),...,\mathbf{y}^{T}(t_{n})]^{T}$ can be written as
[…]
In the training phase of the RC framework, it is customarily assumed that the reservoir is fixed, whereas the readout nodes must be trained. Thus, the optimization will simplify into a linear regression, which requires minimizing the following loss function:
[…]
This objective function, referred to as linear ridge regression Hoerl and Kennard (1970), can readily calibrate the readout nodes through an explicit solution when the ridge coefficient is considered known, i.e.
$\mathbf{\hat{W}}_{out}=(\mathbf{R^{\it{T}}_{\it{n}}}\mathbf{R_{\it{n}}}+\it{\lambda}\mathbb{I})^{-1}\mathbf{R_{\it{n}}^{\it{T}}}\mathbf{Y_{\it{n}}}$ (4)
[…]
@@ II.2 Proposed Learning Protocol
The reservoir is assumed to take noisy and incomplete measured data as the input and return filtered and complete states as the output. FIG. 1 provides an overview of this specific implementation of the reservoir. At first glance, it might be difficult to train such an RC since both noisy and clean data are required in the training phase. However, such data can be obtained via computer simulations wherein the dynamics are considered entirely known for specific chaotic orbits. Another possibility is to employ simpler system dynamics for acquiring noisy data and attempting to clean those measurements through conventional filtering techniques Donoho (1995). In both cases, the RC trained based on these data sets is anticipated to generalize across a large neighborhood such that it can be employed for the denoising and reconstruction of unexplored nonlinear dynamics while being trained on limited simulated chaotic orbits.
Figure 1: Overview of the proposed RC protocol for denoising and reconstruction of nonlinear dynamics. Blue, red, and orange circles, show reservoir nodes, input nodes, and readout nodes, respectively. The blue arrows show the edges of the reservoir chosen randomly. The solid black arrows show the input connection to specific nodes of the reservoir, and the black dashed arrows show the readout connections. (a) Incomplete and noisy dynamics of Lorenz system considered as the input to the RC (b) Reproduced Lorenz dynamics obtained through the RC framework (c) Comparison of the input and output of the reservoir in the phase-space, representing how three-dimensional dynamics (blue curve) are reconstructed from two-dimensional noisy signals (orange curve).
@@ II.3 Hyperparameter Optimization
The RC framework presented in the preceding section is straightforward to follow and implement because it only requires performing a linear regression for the training of the readout nodes. However, the performance of the reservoir computer often benefits from the fine-tuning of the reservoir and optimizing its underlying parameters. Such parameters include the number of neurons, leakage rate, spectral radius, input scaling, and node connectivity, whose optimization can be performed through a greedy algorithm that minimize the loss function of Eq. 3 for the underlying parameters while considering a held-out data set. To formalize such an optimization, let the hyperparameters of the RC be denoted by $\phi$, specifically comprising the number of reservoir nodes, the leakage rate, the spectral radius, the input scaling ratio, and the output connectivity. The optimal values of these hyperparameters should be determined from
$\mathbf{\hat{\phi}}=\arg\min_{\phi}Loss$ (5)
where the loss function is given by Eq. 3. It should be noted that this minimization might encounter computational difficulties if gradient-descent-based optimization approaches Bottou (1999) are employed. In effect, calculating the gradients requires implementing back-propagation Rumelhart, Hinton, and Williams (1986) through the hidden states of the reservoir, which is often deemed challenging, especially when the number of time steps in the training data is large. Therefore, it is recommended to employ surrogate-assisted optimization techniques integrated with cross-validation techniques, which avoid computing the gradients Li, Small, and Lei (2023) as realized in the python-based package scikit-optimizeBergstra and Bengio (2012).

### v2

@@ Abstract
Measurements acquired from distributed physical systems are often sparse and noisy. Therefore, signal processing and system identification tools are required to mitigate noise effects and reconstruct unobserved dynamics from limited sensor data. However, this process is particularly challenging because the fundamental equations governing the dynamics are largely unavailable in practice. Reservoir Computing (RC) techniques have shown promise in efficiently simulating dynamical systems through an unstructured and efficient computation graph comprising a set of neurons with random connectivity. However, the potential of RC to operate in noisy regimes and distinguish noise from the primary smooth or non-smooth deterministic dynamics of the system has not been fully explored. This paper presents a novel RC method for noise filtering and reconstructing unobserved nonlinear dynamics, offering a novel learning protocol associated with hyperparameter optimization. The performance of the RC in terms of noise intensity, noise frequency content, and drastic shifts in dynamical parameters is studied in two illustrative examples involving the nonlinear dynamics of the Lorenz attractor and the adaptive exponential integrate-and-fire system. It is demonstrated that denoising performance improves by truncating redundant nodes and edges of the reservoir, as well as by properly optimizing hyperparameters, such as the leakage rate, spectral radius, input connectivity, and ridge regression parameter. Furthermore, the presented framework shows good generalization behavior when tested for reconstructing unseen and qualitatively different attractors. Compared to the extended Kalman filter, the presented RC framework yields competitive accuracy at low signal-to-noise ratios and high-frequency ranges.
@@
Reconstruction of unobserved dynamical responses from sparse and noisy observations is a crucial and challenging task when dealing with unknown nonlinear systems. Filtering techniques have commonly been employed to handle such problems, but they often demand a thorough understanding of the underlying processes, such as the governing equations in terms of a physics-based state-space model. This paper presents a novel truncated reservoir computing approach to simulate nonlinear dynamics and implicitly recover the underlying physics from noisy measurements. It is proposed that reservoir computers can effectively distinguish noise from dynamics and reconstruct unmeasured dynamical responses from noisy ones without having to involve any supplementary noise reduction techniques. While the RC acts as a model-free mathematical modeling technique, it offers competitive accuracy compared to the conventional filtering techniques at certain noise levels.
@@ I Introduction
Measurements of physical quantities from distributed systems are often noisy and incomplete in covering all response quantities. This issue has inspired the development of a wide range of denoising and reconstruction methods, which fall under two major classes, namely model-free and model-based methods, depending on whether they demand prior knowledge of underlying physics. A good example of the former is wavelet-based denoising Donoho (1995), which does not require incorporating physical knowledge explicitly, whereas stochastic filtering techniques, including Kalman-type filters Anderson and Moore (1979), often demand a general understanding of the physics, characterized via state-space models. However, the foundation of both methods primarily rests upon a key hypothesis, implying that the desired signal and the noise have distinguishable low and high-dimensional features in a specific latent space.
@@
Reservoir computing is a computational framework rooted in the broader domain of recurrent neural networks (RNNs) and dynamical systems theory. It leverages the dynamics of a fixed, randomly initialized high-dimensional system, known as the “reservoir,” to transform sequential input data into a latent space representation. In this paradigm, only the output layer is trained through ridge regression Ghani et al. (2010), which linearly combines the reservoir states to minimize prediction errors. This significantly reduces the computational complexity associated with the training, mainly because only a system of linear equations needs to be solved coupled with a proper regularization. RC has been particularly effective in time-series processing Gauthier et al. (2021); Thorne et al. (2022); Shahi, Fenton, and Cherry (2022a), speech recognition Ghani et al. (2010); Araujo et al. (2020), and chaotic system modeling Pathak et al. (2017); Krishnagopal et al. (2020); Carroll (2022) due to its ability to exploit the reservoir’s rich temporal dynamics and memory capacity. Heuristics of building the reservoir formulation, e.g., including an output feedback or a delayed input, have been shown to improve the results Jaurigue et al. (2021). The generalization behavior of RC has recently been studied, where an RC approach is employed for making predictions on unexplored attractors Flynn, Tsachouridis, and Amann (2021); Parlitz (2024). However, the optimal size and structure of the reservoir remain an open question since reservoir networks are mostly initialized considering random connectivity. Only lately researchers have started investigating the structure-function relationships to discover network properties that are beneficial for a given learning task within the RC framework Yadav, Sinha, and Stender (2025); Yadav et al. (2025).
@@
Recent research has explored the impact of noise on the RC’s performance, revealing its strength and weaknesses for signal reconstruction across various contexts. Rarely RC’s have been employed specifically for simultaneous denoising and signal reconstruction tasks. Semenova et al. Semenova et al. (2019) have studied the propagation of noise from the input nodes toward the reservoir outputs in recurrent and multilayer networks. Estébanez et al. Estébanez, Fischer, and Soriano (2019) have shown that, while noise can degrade the performance of RC-based analog computations, its introduction in the training phase can lead to improved robustness and accuracy. Similarly, Wikner et al. Wikner et al. (2024) have shown that adding noise to the training data can improve the stability of predictions. Röhm et al. Röhm, Gauthier, and Fischer (2021) have shown that an RC trained based on a single noisy trajectory may still be able to reproduce unexplored attractors. Nathe et al. Nathe et al. (2023) have investigated the impact of measurement noise on the performance of RC when used for reconstructing chaotic systems. Lee et al. Lee et al. (2025) have employed RC along with a nonlinear forward operator for efficient image denoising. Choi and Kim Choi and Kim (2024) have separated the signal from noise as the predictable part modeled by RC and acquired information about the additive or multiplicative nature of the noise, as well as noise statistical properties. Despite these advances in RC-based denoising approaches, a thorough investigation is still required to explore the potential of RC for discarding different noise processes and factors affecting denoising performance.
@@
This work introduces a novel learning protocol for denoising and reconstruction of nonlinear dynamics based on a truncated RC framework. This approach begins with training the reservoir on a specific data set using the conventional ridge regression, followed by a new optimization strategy that calibrates key hyperparameters, including the ridge regression coefficient, leakage rate, and spectral radius, aiming to enhance performance. Once the optimized reservoir is obtained, we truncate redundant nodes and edges according to the model accuracy on the validation dataset using a novel pruning strategy. Two numerical examples are provided to demonstrate the proposed protocol and study the impact of noise intensity and frequency content on the performance of the RC. The generalization behavior of the proposed networks is demonstrated against drastic changes in the dynamical parameters and Signal-to-Noise Ratio (SNR). Comparison with the Extended Kalman Filter (EKF), a near optimal filter Anderson and Moore (1979), is provided for the Lorenz attractor example, considering different noise levels. Finally, the method is further demonstrated using the non-smooth deterministic dynamics of the adaptive exponential integrate-and-fire (AdEx) model.
This study primarily investigates the potential of reservoir computing for denoising and state reconstruction. In addition, it applies hyperparameter optimization and explores the pruning of redundant nodes and edges. The pruning of small reservoir networks, specifically the removal of redundant nodes and connections, has received relatively little attention in previous studies. Notably, Dutoit et al Dutoit et al. (2009) have shown that the pruning of the readout nodes is an effective strategy to improve the prediction accuracy and reduce hardware and software computational demands. Scardapane et al. Scardapane et al. (2015) have characterized the significance of reservoir synapses according to the correlation between the input and output of the neurons and specified the significance of neurons based on the importance of their neighboring synapses. Then, based on this criterion, they have presented a pruning approach for reducing the reservoir dimensionality without calculating the error gradients Scardapane et al. (2015). Haluszczynski et al. Haluszczynski et al. (2020) have shown that removing the nodes with negative impact on the reservoir’s response to the input can enhance the accuracy and reduce instabilities. Compared to these works, this study provides a more systematic and generic approach to reducing the reservoir nodes and edges.
@@
The remainder of this paper is organized as follows: first, the mathematical foundation of the RC framework will be reviewed, presenting the conventional ridge regression method. Then, the reservoir’s hyperparameters will be identified, for which an optimization approach will be presented. Truncation of the RC through a pruning approach will be discussed, and further comparisons will be made between the proposed RC framework and conventional filtering techniques from a theoretical standpoint. The presented approaches will be demonstrated using two nonlinear systems, aiming to address simultaneous state reconstruction and denoising tasks. Finally, further discussions and conclusions will be drawn about the denoising performance of the proposed truncated RC framework.
@@ II.1 Echo State Network
Several neural network architectures are available for constructing reservoir computers, among which Echo State Network (ESN) Jaeger (2007) and Liquid State Machines Maass, Natschlaeger, and Markram (2002); Hazan and Manevitz (2012) are perhaps the most widely used ones. In this paper, ESN is employed for creating a reservoir whose formulation is described in terms of the latent dynamical states:
$\mathbf{r}(t_{i+1})=(1-{\it{\alpha}})\mathbf{r}(t_{i})+{\it{\alpha}}\ f(\mathbf{W}_{\mathrm{res}}\mathbf{r}(t_{i})+\mathbf{W}_{\mathrm{in}}\mathbf{u}(t_{i})+\mathbf{b})$ (1)
where $f(.)$ is the activation function, considered $\tanh(.)$, $\mathbf{r}(t_{i})$ is the reservoir state at time $t_{i}$, $\alpha$ is the leakage rate representing how much of the preceding state contributes to the present state, $\mathbf{W}_{\mathrm{res}}$ is the reservoir weight matrix, considered a random weighted Erdős–Rényi (ER) graph matrix with $p$ internal connectivity probability, $\mathbf{W}_{\mathrm{in}}$ is the input weight matrix scaled by a factor of $\zeta$, $\mathbf{u}(t_{i})$ is the input vector at time $t_{i}$, and $\mathbf{b}$ is the bias term. Notably, an important parameter of the reservoir matrix ($\mathbf{W}_{\mathrm{res}}$) is its spectral radius, denoted by $\gamma$ and calculated as the greatest eigenvalue of the reservoir matrix Ren and Ma (2022).
Given the above equation, a recursive formulation is available for calculating the reservoir states. By collecting all states into a matrix like $\mathbf{R}_{n}=[\mathbf{r}(t_{1}),...,\mathbf{r}(t_{n})]^{T}$, simulated responses of the physical system at different time steps comprised in $\mathbb{Y}_{n}=[\mathbf{y}(t_{1}),...,\mathbf{y}(t_{n})]^{T}$ can be described as
$\mathbf{Y_{\it{n}}}=\mathbf{R_{\it{n}}}\mathbf{W}_{\mathrm{out}}$ (2)
In the training phase of the RC framework, it is customarily assumed that the reservoir is fixed, whereas the readout matrix $\mathbf{W}_{\mathrm{out}}$ must be trained. Thus, the optimization will simplify into a linear regression, which requires minimizing the following loss function:
$L=\|\mathbf{{\hat{Y}_{\it{n}}}-\mathbf{R_{\it{n}}}\mathbf{W}_{\mathrm{out}}}\|_{2}^{2}+\lambda\|\mathbf{W}_{\mathrm{out}}\|_{2}^{2}$ (3)
where $\mathbf{\hat{Y}_{\it{n}}}$ is a vector of actual measurements fed into the RC as input. This formulation, referred to as linear ridge regression Hoerl and Kennard (1970), can readily calibrate the readout nodes through an explicit solution, i.e.,
$\mathbf{\hat{W}}_{\mathrm{out}}=(\mathbf{R^{\it{T}}_{\it{n}}}\mathbf{R_{\it{n}}}+\it{\lambda}\mathbb{I})^{-1}\mathbf{R_{\it{n}}^{\it{T}}}\mathbf{\hat{Y}_{\it{n}}}$ (4)
where ${\hat{\mathbf{W}}}_{\mathrm{out}}$ is the optimal readout matrix, and $\lambda$ is the ridge coefficient, often selected to be a small coefficient like $10^{-8}$. In Eq. 3, it is evident that the first term, $\|{{{\mathbf{\hat{Y}_{\it{n}}}}}-\mathbf{R_{\it{n}}}\mathbf{W}_{\mathrm{out}}}\|_{2}^{2}$, provides a measure of fitting accuracy, whereas the second term, $\lambda\|\mathbf{W}_{\mathrm{out}}\|_{2}^{2}$, involves a penalty over the non-zero elements of $\mathbf{W}_{\mathrm{out}}$, implementing a smooth sparsity. Note that in problems involving different levels of noise, $\lambda$ must be inferred properly based on the data.
@@ II.2 Proposed Learning Protocol
The reservoir is assumed to take noisy and incomplete measured data $\mathbf{U_{\it{n}}}=[\mathbf{u}(t_{1}),...,\mathbf{u}(t_{n})]^{T}$ as the input and return filtered and complete states ${\mathbf{Y_{\it{n}}}}$ as the output. FIG. 1 provides an overview of this specific implementation of the reservoir. At first glance, it might be difficult to train such an RC since both noisy and clean data are required in the training phase. However, such data can be obtained via computer simulations wherein the dynamics are considered entirely known for specific chaotic orbits. Another possibility is to employ simpler system dynamics for acquiring noisy data and attempting to clean those measurements through conventional filtering techniques Donoho (1995). In both cases, the RC trained based on these data sets is anticipated to generalize across a large neighborhood such that it can be employed for the denoising and reconstruction of unexplored nonlinear dynamics while being trained on limited clean data.
Figure 1: Overview of the proposed RC protocol for denoising and reconstruction of nonlinear dynamics. Blue, red, and orange circles show reservoir nodes, input nodes, and readout nodes, respectively. The blue arrows show the edges of the reservoir chosen randomly. The solid black arrows show the input connection to specific nodes of the reservoir, and the black dashed arrows show the readout connections. (a) Incomplete and noisy dynamics of the Lorenz system considered as the input to the RC (b) Reproduced Lorenz dynamics obtained through the RC framework (c) Comparison of the input and output of the reservoir in the phase space, representing how three-dimensional dynamics (blue curve) are reconstructed from two-dimensional noisy signals (orange curve).
@@ II.3 Hyperparameter Optimization
The RC framework presented in the preceding section is straightforward to follow and implement because it only requires performing a linear regression for the training of the readout unit. However, the performance of the reservoir computer often benefits from the fine-tuning of the reservoir and optimizing its underlying parameters. Such parameters include the number of neurons, leakage rate, spectral radius, input scaling, and node connectivity, whose optimization can be performed through a greedy algorithm that minimizes the loss function of Eq. 3 for the underlying parameters while considering a held-out data set. To formalize such an optimization, let the hyperparameters of the RC be denoted by $\phi$, specifically comprising the number of reservoir nodes ($N$), the leakage rate ($\alpha$), the spectral radius ($\gamma$), the input scaling ratio ($\zeta$), and the reservoir internal connectivity rate ($p$). The optimal values of these hyperparameters should be determined from
$\mathbf{\hat{\phi}}=\arg\min_{\phi}\left(L\right)$ (5)
where the loss function is given by Eq. 3. It should be noted that this minimization might encounter computational difficulties if gradient-descent-based optimization approaches Bottou (1999) are employed. In effect, calculating the gradients requires implementing back-propagation Rumelhart, Hinton, and Williams (1986) through the hidden states of the reservoir, which is often deemed challenging if not impossible, especially when the number of time steps in the training data is large Hanin (2018). Therefore, grid-based Shahi, Fenton, and Cherry (2022b) or random search algorithms Kaelo and Ali (2006) are often preferred in the literature. In this study, surrogate-assisted optimization techniques are employed along with cross-validation algorithms, which avoid computing the gradients Li, Small, and Lei (2023) using the Python-based package scikit-optimizeBergstra and Bengio (2012).